# FMDA-OT: Federated Multi-source Domain Adaptation Through Optimal Transport

### v1

@@ Abstract
Multi-source Domain Adaptation (MDA) aims to adapt models trained on multiple labeled source domains to an unlabeled target domain. In this paper, we introduce our approach as a collaborative MDA framework, which comprises two adaptation phases. Firstly, we conduct domain adaptation for each source individually with the target, utilizing optimal transport. Then, in the second phase, which constitutes the final part of the framework, we design the architecture of centralized federated learning to collaborate the N models representing the N sources. This architecture offers the advantage of using the sources without accessing their data, thus resolving data privacy issues inherent in domain adaptation. Additionally, during this phase, the server guides and fine-tunes the adaptation using a small number of pseudo-labeled samples available in the target domain, referred to as the target validation subset of the dataset.
@@ 1 Introduction
Supervised learning is a pillar of machine learning, thus this supervision requires labels, which is most of the time an expensive task. While transfer learning and generalisation can be seen as solutions, but due to domain shift, in other words, divergence between the domains (i.e. the difference between probability distributions of two domains), the task model trained on labelled data from domain A (source) and tested on unlabelled data from domain B (target) will have a decay in the performance in comparison with training and testing on the same domain B, which is out capacity in the absence of labels. To address these challenges, domain adaptation approaches are employed. Within these approaches, some can manage adaptation between a single source and a single target, referred to as single-source domain adaptation (SDA) methods, while others can handle adaptation between multiple sources and a single target, known as Multi-source domain adaptation (MDA) methods.
@@
Several works have been proposed in MDA, [Mansour et al., 2021] defended a theoretical analysis considering the assumption says that the target is a convex combination of source domains. [Zhao et al., 2019] Proposed new error bounds for both regression and classification tasks. There are other works proposed on specific algorithms. [Zhao et al., 2020] Combines diverse classifiers by a novel weighting technique. [Hoffman et al., 2012] Introduced a domain transform framework that finds the latent space of the domains by using a clustering approach. These researches generally, either estimate a mixed distribution or combine with a specific approach multiple single-source models.
Collaborative learning paradigm generally and Federated Learning (FL) particularly has attracted the attention of the academia and industry communities. The term Federated Learning was introduced by [McMahan et al., 2017]. This new machine learning approach permits to train a model on different data from different sources that can be mobile devices or even organizations. Our implemented version of federated learning will be used in our domain adaption approach, where each source domain will be considered and represented as a data source in the federated learning paradigm. This approach will consider and open the door for data privacy in domain adaptation where we do not have the access to the source domains data.
@@
Unsupervised domain adaptation methods relying on optimal transport have gained attention due to the recent success of optimal transport in diverse machine learning problems [Laclau et al., 2017] [Ben-Bouazza et al., 2019] [Ben-Bouazza et al., 2022]. These domain adaptation models have increasingly captured interest thanks to their ability to explicitly discover and learn the lowest transformation cost from the source to the target domains. This capability enables them to effectively reduce the discrepancy between the two domains.
Addressing this limitations we propose our approach which combines between optimal transport and federated learning to train a model on multiple source domains and have better performance on the target domain. This paper will be organised in the rest as: Section 2 offers the required preliminary and preparatory knowledge beside notations for unsupervised multi-source domain adaptation, as well as optimal transport theory for domain adaptation. Section 3 has been devoted to present the proposed approach for MDA, which is named Fedetared Multi-source Domain Adaptation through Optimal Transport (FMDA-OT). Section 4 demonstrates the performed comparative study with state-of-the-art methods on two benchmark datasets for MDA. To conclude in section 5.
@@ 2.1 Unsupervised Multi-Source Domain Adaptation
Unsupervised Multi-source domain adaptation (UMDA) is a transductive learning problem, therefore, both labelled data from the source domains and unlabelled data from the target domain are assumed to be accessible throughout the training phase [Redko et al., 2019].
More formally, for each source, we have access to a set $S_{i}=\{(x_{j}^{i},y_{j}^{i})\}_{j=1}^{n_{i}}$ of $n_{i}$ independent and identically distributed (i.i.d.) labeled samples of the $i^{th}$ source pulled from the joint distribution $\mathcal{S}_{i}$, and a set $T=\{x_{j}\}_{j=1}^{m}$ of $m$ i.i.d. unlabeled target samples pulled from the marginal distribution $\mathcal{T}_{\mathcal{X}}$, of the joint distribution $\mathcal{T}$ over $\mathcal{X}$, properly:
@@
The objective of unsupervised multi-source domain adaptation is to build a robust classifier $\eta:\mathcal{X}\to\mathcal{Y}$ with a slight target risk:
@@ 2.2 Optimal Transport
$\mu=\sum_{i=1}^{n}\alpha_{i}\delta_{x_{i}}\,\,\text{and}\,\,\nu=\sum_{j=1}^{m}\beta_{j}\delta_{y_{j}}\vspace{3mm}$ (1)
@@
The discrete Kantorovich problem is a linear program and can therefore be solved exactly in $\mathcal{O}(r^{3}log(r))$, where $r=\max(n,m)$ using simplex algorithm and interior point methods, that is a heavy calculation price. In practice, An entropy smoothing variant [Cuturi, 2013] leads to much more efficient algorithms:
@@
$(\mathcal{MK}^{\,\varepsilon})\,\,\,\,\,\,\underset{\gamma\in U(\alpha,\beta)}{\min}\langle{\gamma},{C}\rangle_{F}-\varepsilon\mathcal{H}(\gamma),\vspace{2mm}$ (5)
@@
where $\mathcal{H}(\gamma)=-\sum_{i=1}^{n}\sum_{j=1}^{m}\gamma_{ij}(\log(\gamma_{ij})-1)$ is the entropy of $\gamma$. The solution to this strictly convex optimization problem is of the form $\gamma^{*}=diag(u)Kdiag(v)$, with $K=\exp(\frac{-C}{\varepsilon})$, and can be obtained efficiently via the Sinkhorn-Knopp algorithm, an iterative matrix scaling method [Peyré et al., 2019].
2.3 Federated Learning
@@
Federated learning enables us to jointly train a machine learning model through the collaboration of multiple parties (i.e. devices, organisations …etc.) without exchanging the local data. Generally federated learning approaches can be classified into either centralised or decentralised paradigm. Many federated learning systems today support neural network training. Google proposed a scalable production system that authorizes the training of deep neural networks by hundreds of millions of devices [Bonawitz et al., 2019]. Centralised FL can be represented as a set of N models for a set of N clients (i.e data sources) and a server model that centralises the learning process. Figure 1 shows the centralised FL paradigm.
Figure 1: Centralised Federated Learning
@@
The training process in centralized federated learning (FL) can be viewed as consisting of four repetitive phases. Firstly, each client model is trained on its local private data. Then, as a second phase, each client sends its model (i.e., weights) to the server following the FedAvg algorithm [McMahan et al., 2017]. Various algorithms have been proposed to handle the exchange of information between models, one of which is the FedSGD algorithm [McMahan et al., 2017]. This approach allows for the transmission of gradients rather than weights, which is considered more effective but requires increased communication, potentially straining the distributed FL system. The third phase involves the server averaging the received weights from the N clients. This average is weighted by the number of samples in each client’s data, normalized by the sum of the counts of all samples across all clients, as proposed by Google [McMahan et al., 2017]. Finally, in the fourth phase, the clients receive the new weights and update their models.
@@
The decentralized or Peer-to-Peer federated learning paradigm follows the same principles as centralized FL, except it doesn’t require a trusted server to centralize the learning process. Figure 2 illustrates the overall repetitive process of training N client models.
Figure 2: Decentralised Federated Learning
@@
Initially, each client trains its model on its local private data. Then, in the second and third phases, it requests the weights or gradients from other clients to update its model with its local data and the received exchangeables (i.e., weights or gradients). Many algorithms have been proposed in this paradigm; BrainTorrent [Roy et al., 2019] is a selective algorithm in which clients communicate with each other and update their models using only the exchangeables from clients who have updates to send. This process continues until convergence. It’s evident that the order of operations is crucial in decentralized federated learning algorithms and significantly impacts the final training results. The outcome may differ if, for instance, client 1 uses the weights or gradients from client 2 before client 2 does. This challenge emphasizes the importance of collaboration order. simFL [Li et al., 2020] is a decentralization of federated learning that adopts the centralized paradigm in a decentralized manner. The proposal involves selecting one client as the server in each iteration and treating the others as clients. The model is then updated based on the chosen client’s weights and the received ones, before being sent back to the rest. However, simFL may encounter security issues or instability due to the variable server.
@@ 3 Our Approach
3.1 Overall framework of FMDA-OT
Figure 3 illustrates the overall framework of our proposition, which can be divided into two main parts of the adaptation. Initially, each client from the source domains receives a small amount of target domain data, which will be pseudo-labeled for reasons discussed later in the paper. In this phase, we apply the optimal transport to project each source domain data into the target domain space using the Sinkhorn algorithm with L1L2 class regularization. This process aims to find a new representation of the source domain data that closely resembles the representation of the target domain, thereby reducing the domain shift phenomenon between them. This new representation is referred to as the transported data of the source domain.
Figure 3: Overall Framework of FMDA-OT
Optimal transport generates new representations of the source domain data, which may not always be the most discriminative representation. In some cases, it could be inferior to the original representation of the source domain data, possibly due to factors such as data imbalance. To address this issue, we utilize pseudo-labeled data from the target domain to assess the quality of the new representation. This involves testing and comparing the accuracy of models trained on both the transported and original data. If an improvement is observed when testing on the transported data, the client may opt to use this data for further step of the adaptation. Otherwise, the original representation is retained as the preferred choice. The subsequent step involves the intervention of centralized federated learning, with each client acting as a source domain. In our approach, we employ the FedAvg algorithm for the learning process. Here, each client sends its own weights after the local training step, and the server aggregates these weights, taking into account coefficients assigned to each client.
Google, in its proposal [McMahan et al., 2017], defines the weight aggregation that takes place on the server as a weighted sum of all the clients’ weights. The coefficient assigned to each client is directly proportional to the number of samples in that client. Formally:
@@ 3.1 Overall framework of FMDA-OT
While it is logical and sensible to weight the sum of the weights based on the number of samples provided by each client, this approach appears to work better in scenarios where all clients have the same probability distribution (i.e., the same domain). In unsupervised Multi-source Domain Adaptation (MDA), where different domains are involved, a client with fewer samples may contribute more effectively to improving the final model than another client with more data. Failing to account for this domain shift and weighting solely based on the number of samples can lead to less reliable results. To address this issue, we propose weighting all clients by defining coefficients that are directly proportional to the performance of each client model when tested on the pseudo-labeled portion of target domain data.
Our proposition takes into account the paramount importance of privacy, which is a necessity across all industries today. It is often challenging to train a model without direct access to the data, particularly in the context of domain adaptation and phenomena such as domain shift, where data from multiple sources may be required to leverage information for building a robust model while still adhering to privacy regulations. To enhance domain adaptation within the constraints of privacy, we propose FMDA-OT. The following Table 1 summarizes the advantages introduced by FMDA-OT.
@@
FMDA-OT (Ours)
3.2 Pseudo labelling the validation data of target domain
The significance of this validation proportion lies in its ability to guide adaptation, thereby reducing domain shift by maintaining a dynamic training process. In our scalable proposition, this validation proportion can be labeled by experts if available, which can lead to improved results. Alternatively, we propose a pseudo-labeling approach to handle full unsupervised training for the target domain model. The HOT [El Hamri et al., 2022] approach offers a scalable solution that can be applied in our scenario. Initially, we cluster the target domain data on the server using one of the proposed clustering algorithms. In our experiments, we utilized spectral clustering, as it represents the state-of-the-art in clustering algorithms. Once the clusters are obtained and random labels are assigned, this pseudo-labeled validation data is distributed to each client. Here, we apply hierarchical optimal transport (HOT), which calculates the correspondence between each source domain label and the random labels assigned to this data during clustering. Figure 4 illustrates the HOT correspondence matrix.
@@ 3.2 Pseudo labelling the validation data of target domain
Not necessarily do we obtain identical results from one client to another. For this reason, if the number of clients is three or more, we apply a majority vote. In cases where we have only two clients, we consider the correspondence of the closest source domain to the target according to the Wasserstein distance.
3.3 On-clients Testing And Server Weights Aggregation
The validation part of the target data serves as a guide for adaptation in both phases. During the optimal transport phase, this part determines whether accuracy improves when using the new representation obtained by the optimal transport transformation. If not, the original data representation is retained. This testing occurs across all clients with all sources. Each client receives this pseudo-labeled data and conducts an internal test to determine which data to use in the next phase of adaptation, which is federated learning.
In the federated learning training phase, the server utilizes this part. Upon receiving weights from all clients, the server tests the received model from each client on this validation part and records the accuracy. This process is repeated for all client weights. The accuracy of all clients in the federated learning is then normalized by their sum and used as weights for aggregating all incoming weights. This approach maintains highly dynamic adaptation, adjusting the model with changes occurring in each client model during the learning process. It also provides a robust metric for the importance of each source to the target.
Besides privacy, federated learning’s collaborative aspect is a significant advantage. Models share weights, addressing each model’s shortcomings. Through collaboration, we can harness the full potential of data and knowledge without sharing actual data or compromising privacy.
The guarantee of this pseudo-labeling is directly related to clustering quality. Higher cluster purity leads to better label assignments. Hence, we used spectral clustering in experiments. HOT [El Hamri et al., 2022] effectively projects random labels from clustering to real ones. This pseudo-labeling keeps the training process entirely unsupervised, yet flexible enough to replace the validation part with expert-labeled data for better weight fine-tuning. In cases of pure clusters, pseudo labels match real ones perfectly.

### v2

# Collaborative Multi-source Domain Adaptation Through Optimal Transport

@@ Abstract
Multi-source Domain Adaptation (MDA) seeks to adapt models trained on data from multiple labeled source domains to perform effectively on an unlabeled target domain data, assuming access to sources data. To address the challenges of model adaptation and data privacy, we introduce Collaborative MDA Through Optimal Transport (CMDA-OT), a novel framework consisting of two key phases. In the first phase, each source domain is independently adapted to the target domain using optimal transport methods. In the second phase, a centralized collaborative learning architecture is employed, which aggregates the N models from the N sources without accessing their data, thereby safeguarding privacy. During this process, the server leverages a small set of pseudo-labeled samples from the target domain, known as the target validation subset, to refine and guide the adaptation. This dual-phase approach not only improves model performance on the target domain but also addresses vital privacy challenges inherent in domain adaptation.
@@ 1 Introduction
Supervised learning is a cornerstone of machine learning, but it relies heavily on labeled data, which is often expensive and time-consuming to obtain. While transfer learning and generalization offer potential solutions, domain shift—defined as the divergence between the probability distributions of two domains—remains a significant challenge. Consequently, a model trained on labeled data from domain A (source) and tested on unlabeled data from domain B (target) typically suffers from performance degradation compared to being trained and tested within the same domain B, which is impractical without labels. To address these challenges, domain adaptation methods are employed. Within these approaches, some can manage adaptation between a single source and a single target, referred to as single-source domain adaptation (SDA) methods, while others handle adaptation between multiple sources and a single target, known as multi-source domain adaptation (MDA) methods.
@@
Several works have advanced MDA. For instance, [Mansour et al., 2021] provided a theoretical analysis based on the assumption that the target is a convex combination of source domains. [Zhao et al., 2019] proposed new error bounds for regression and classification tasks. Other research efforts focused on specific algorithms, such as combining diverse classifiers through novel weighting techniques [Zhao et al., 2020], or using domain transform frameworks to find latent domain spaces via clustering [Hoffman et al., 2012]. Generally, these studies estimate a mixed distribution or combine multiple single-source models.
The collaborative learning paradigm generally, and federated learning particularly, has garnered significant attention from both academia and industry. Initially introduced by [McMahan et al., 2017], this approach enables model training on data from various sources, such as mobile devices or organizations, without requiring centralized data access. This concept is pivotal for addressing data privacy concerns in domain adaptation, as it allows for the utilization of source domain data without direct access.
@@
Unsupervised domain adaptation methods based on optimal transport have recently gained traction due to their success in various machine learning problems [Laclau et al., 2017] [Ben-Bouazza et al., 2019] [Ben-Bouazza et al., 2022]. These methods are adept at discovering and learning the minimal transformation cost from source to target domains, effectively reducing domain discrepancies (i.e., domain shift).
In this context, we propose a novel approach that combines optimal transport with collaborative learning to enhance model performance on the target domain using multiple source domains. Our method, which we term Collaborative Multi-source Domain Adaptation through Optimal Transport (CMDA-OT), leverages the strengths of both frameworks to achieve superior adaptation while preserving data privacy.
The rest of this paper is structured as follows: Section 2 provides the necessary preliminary knowledge and notations for unsupervised multi-source domain adaptation and optimal transport theory. Section 3 presents our proposed approach for MDA, CMDA-OT. Section 4 demonstrates a comparative study with state-of-the-art methods on two benchmark datasets for MDA. Finally, Section 5 concludes the paper.
@@ 2.1 Unsupervised Multi-Source Domain Adaptation
Unsupervised Multi-source Domain Adaptation (UMDA) is a transductive learning problem. Both labeled data from the source domains and unlabeled data from the target domain are assumed to be accessible throughout the training phase [Redko et al., 2019].
More formally, for each source, we have access to a set $S_{i}=\{(x_{j}^{i},y_{j}^{i})\}_{j=1}^{n_{i}}$ of $n_{i}$ independent and identically distributed (i.i.d.) labeled samples of the $i^{th}$ source pulled from the joint distribution $\mathcal{S}_{i}$, and a set $T=\{x_{j}\}_{j=1}^{m}$ of $m$ i.i.d. unlabeled target samples pulled from the marginal distribution $\mathcal{T}_{\mathcal{X}}$ of the joint distribution $\mathcal{T}$ over $\mathcal{X}$:
@@
The objective of unsupervised multi-source domain adaptation is to build a robust classifier $\eta:\mathcal{X}\to\mathcal{Y}$ with a low target risk:
@@ 2.2 Optimal Transport
$\mu=\sum_{i=1}^{n}\alpha_{i}\delta_{x_{i}}\,\,\text{and}\,\,\nu=\sum_{j=1}^{m}\beta_{j}\delta_{y_{j}}$ (1)
@@
The discrete Kantorovich problem is a linear program and can therefore be solved exactly in $\mathcal{O}(r^{3}\log(r))$, where $r=\max(n,m)$ using the simplex algorithm and interior point methods, which is computationally intensive. In practice, an entropy smoothing variant [Cuturi, 2013] leads to much more efficient algorithms:
@@
$(\mathcal{MK}^{\,\varepsilon})\,\,\,\,\,\,\underset{\gamma\in U(\alpha,\beta)}{\min}\langle{\gamma},{C}\rangle_{F}-\varepsilon\mathcal{H}(\gamma),$ (5)
@@
where $\mathcal{H}(\gamma)=-\sum_{i=1}^{n}\sum_{j=1}^{m}\gamma_{ij}(\log(\gamma_{ij})-1)$ is the entropy of $\gamma$. The solution to this strictly convex optimization problem is of the form $\gamma^{*}=\text{diag}(u)K\text{diag}(v)$, with $K=\exp(\frac{-C}{\varepsilon})$, and can be obtained efficiently via the Sinkhorn-Knopp algorithm, an iterative matrix scaling method [Peyré et al., 2019].
2.3 Collaborative Learning with Federated Learning as Instance
@@
Collaborative learning enables the joint training of machine learning models through the collaboration of multiple parties (e.g., devices, organizations) without sharing local data. Federated learning (FL) is a specific instance of collaborative learning and is typically classified into either centralized or decentralized paradigms. Many contemporary FL systems support neural network training. Google proposed a scalable production system that allows the training of deep neural networks by hundreds of millions of devices [Bonawitz et al., 2019]. Centralized FL can be represented by a set of $N$ models for $N$ clients (i.e., data sources) and a server model that centralizes the learning process, as depicted in Figure 1.
Figure 1: Centralized Federated Learning
@@
The training process in centralized FL consists of four repetitive phases. Firstly, each client model is trained on its local private data. Secondly, each client sends its model (i.e., weights) to the server, following the FedAvg algorithm [McMahan et al., 2017]. Various algorithms, such as the FedSGD algorithm [McMahan et al., 2017], have been proposed to handle the exchange of information between models. The FedSGD algorithm allows for the transmission of gradients rather than weights, which is considered more effective but requires increased communication, potentially straining the distributed FL system. In the third phase, the server averages the received weights from the $N$ clients. This average is weighted by the number of samples in each client’s data, normalized by the sum of the counts of all samples across all clients, as proposed by Google [McMahan et al., 2017]. Finally, in the fourth phase, the clients receive the new weights and update their models.
@@
The decentralized or Peer-to-Peer federated learning paradigm follows the same principles as centralized FL, except it does not require a trusted server to centralize the learning process. Figure 2 illustrates the overall repetitive process of training $N$ client models.
Figure 2: Decentralized Federated Learning
@@
Initially, each client trains its model on its local private data. In the second and third phases, it requests the weights or gradients from other clients to update its model with both its local data and the received exchangeables (i.e., weights or gradients). Numerous algorithms have been proposed within this paradigm. For instance, BrainTorrent [Roy et al., 2019] is a selective algorithm where clients communicate and update their models using only the exchangeables from clients who have updates to send. This process continues until convergence. The order of operations is crucial in decentralized FL algorithms and significantly impacts the final training results. For example, the outcome may differ if client 1 uses the weights or gradients from client 2 before client 2 does. This challenge underscores the importance of the collaboration order. simFL [Li et al., 2020] is a decentralized approach to federated learning that adopts the centralized paradigm in a decentralized manner. This method involves selecting one client as the server in each iteration and treating the others as clients. The model is updated based on the chosen client’s weights and the received ones before being sent back to the rest. However, simFL may encounter security issues or instability due to the variable server.
@@ 3 Our Approach
3.1 Overall Framework of CMDA-OT
Figure 3 illustrates the overall collaborative framework of our proposition, which can be divided into two main parts of the adaptation. Initially, we apply optimal transport, as the first part, to project each source domain data into the target domain space using the Sinkhorn algorithm with L1L2 class regularization. This process aims to find a new representation of the source domain data that closely resembles the representation of the target domain, thereby reducing the domain shift phenomenon between them. This new representation is referred to as the transported data of the source domain.
Figure 3: Overall Framework of CMDA-OT
Optimal transport generates new representations of the source domain data, which may not always be the most discriminative representation. In some cases, it could be inferior to the original representation of the source domain data, possibly due to factors such as data imbalance. To address this issue, we utilize a small set of pseudo-labeled data from the target domain, referred to as the target domain validation data, to assess the quality of the new representation. This involves testing and comparing the accuracy of models trained on both the transported and original data. If an improvement is observed when testing on the transported data, the client may opt to use this data for further steps of the adaptation. Otherwise, the original representation is retained as the preferred choice.
The subsequent part involves the intervention of centralized collaborative learning, federated learning as instance, with each client acting as a source domain. In our approach, we employ the FedAvg algorithm [McMahan et al., 2017] for the learning process. Here, each client sends its own weights after the local training step, and the server aggregates these weights, taking into account coefficients assigned to each client.
[McMahan et al., 2017] defines the weight aggregation that takes place on the server as a weighted sum of all the clients’ weights. The coefficient assigned to each client is directly proportional to the number of samples in that client. Formally:
@@ 3.1 Overall Framework of CMDA-OT
While it is logical and sensible to weight the sum of the weights based on the number of samples provided by each client, this approach appears to work better in scenarios where all clients have the same probability distribution (i.e., the same domain). In Unsupervised Multi-source Domain Adaptation (UMDA), where different domains are involved, a client with fewer samples may contribute more effectively to improving the final model than another client with more data. Failing to account for this domain shift and weighting solely based on the number of samples can lead to less reliable results. To address this issue, we propose weighting all clients by defining coefficients that are directly proportional to the performance of each client model when tested on the small pseudo-labeled validation portion of target domain data. This test takes place in each client.
Our proposition takes into account the paramount importance of privacy, which is a necessity across all industries today. It is often challenging to train a model without direct access to the data, particularly in the context of domain adaptation and phenomena such as domain shift, where data from multiple sources may be required to leverage information for building a robust model while still adhering to privacy regulations. To enhance domain adaptation within the constraints of privacy, we propose CMDA-OT. Table 1 highlights the advantages of CMDA-OT in comparison to relevant literature and related works.
@@
CMDA-OT (Ours)
3.2 Pseudo-Labeling the Target Domain Validation Data
The significance of this validation proportion lies in its ability to guide adaptation, thereby reducing domain shift by maintaining a dynamic training process. In our scalable proposition, this validation proportion can be labeled by experts if available, which can lead to improved results. Alternatively, we propose a pseudo-labeling approach to handle fully unsupervised training for the target domain model. The HOT [El Hamri et al., 2022] approach offers a scalable solution that can be applied in our scenario. Initially, we cluster the target domain data on the server using one of the clustering algorithms proposed by the literature. In our experiments, we utilized spectral clustering, as it represents one of the state-of-the-art in clustering algorithms. Once the clusters are obtained and random labels are assigned, this pseudo-labeled validation data is distributed to each client. Here, we apply hierarchical optimal transport (HOT), which calculates the correspondence between each source domain labels and the random labels assigned to this data during clustering. Figure 4 illustrates an example of the HOT correspondence matrix.
@@ 3.2 Pseudo-Labeling the Target Domain Validation Data
The results obtained from different clients are not always identical. Therefore, when there are three or more clients, we apply a majority vote to determine the outcome. In cases where there are only two clients, we use the correspondence of the source domain that is closest to the target domain, as determined by the Wasserstein distance.
3.3 On-Client Testing and Server Weights Aggregation
The validation part of the target data serves as a guide for adaptation in both phases. During the optimal transport phase, this part determines whether accuracy improves when using the new representation obtained by the optimal transport transformation. If not, the original data representation is retained. This testing occurs across all clients, i.e., all sources. Each client receives this pseudo-labeled data and conducts an internal test to determine which representation, i.e., original or transported, to use in the next phase of adaptation, which is the collaborative learning.
During the collaborative learning training phase, the server utilizes the target domain validation portion. Upon receiving the models weights from all clients, the server tests each client’s model on this validation set and records their accuracies. This process is repeated for all client models. The accuracies are then normalized by their sum and used as coefficients for aggregating the incoming models weights. This approach ensures a highly dynamic adaptation, continuously adjusting the target model as each client’s model evolves during the learning process. It also provides a robust metric for determining the importance of each source to the target domain.
The effectiveness of validation data pseudo-labeling is directly linked to the quality of clustering. Higher cluster purity results in more accurate label assignments. Therefore, we employed spectral clustering in our experiments. The HOT [El Hamri et al., 2022] method effectively maps random labels from clustering to real labels. This approach ensures that the training process remains entirely unsupervised for better fine-tuning the target model weights, while being flexible enough to replace the expert-labeled validation set. When clusters are pure, the pseudo labels closely match the actual labels.
In addition to preserving privacy, collaborative learning provides substantial benefits in scalability and flexibility. By sharing the weights of each source model, we overcome the limitations of individual models when aggregating them into a unified target model. This collaborative approach allows us to fully leverage the potential of data and knowledge without sharing or exposing the actual data or compromising privacy.
3.4 Collaborative Learning Framework : Adaptability and Scalability
This framework leverages the principles of optimal transport and collaborative learning, and although we have implemented it using centralized federated learning as an instance, the framework is designed to be adaptable to various collaborative learning paradigms. In the centralized federated learning paradigm, our approach ensures that we do not have direct access to the client data (i.e., source domain data), thus preserving client privacy. Instead, the federated server manages the learning process and has access to the target data. This implementation demonstrates the versatility of our framework, allowing it to be extended to other collaborative learning settings while maintaining robust privacy-preserving mechanisms.